• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name zhangyi

    Log entry time 16:37:08 on February22,2009

    Entry number 262685

    keyword=Day shift summary

    Day shift summary
    SL : Yi Zhang
    TO: Zhihong Ye
    3rd: Whitney Armstrong

    08:40 HV of BB wire chamber tripped due to dirty beam restart a new run on BB
    08:49 Stop taking data and finish a NMR measurement.
    10:20 LHRS coda died, dead time is 100% and did not get any event for 15 min. restart coda, the new run number is 20330
    10:55 change target to Ref. cell N2(100psi) and change the Kine. of L HRS to 0.9GeV/C, ask for 5uA beam, start a N2 run on both arm
    11:30 Move target to 3He and start production.
    12:49 Lost beam. MCC said beam would come back 1 hour later.
    15:37 MCC get beam back. Ask for a harp scan first.

    BB: 1692, 1693(BB HV trip), 1694, 1695, 1696(N2), 1697, 1698(stop for beam trip.)

    kine. F(1.13GeV/C): 20327, 20328, 20329(bad), 20330
    kine. D(0.9GeV/C): 20331(N2), 20332, 20333(stop for beam trip)